Fill in the parantheses to make statement correct.
-9-(-2.3)= -9+(?)

Answer by jim_thompson5910(35256)

Subtracting a negative is the same as adding. So in this case,

-9-(-2.3)= -9+(2.3)

or

-9-(-2.3)= -9+2.3